# Epistemic Boundaries and Quantum Indeterminacy: What Local Observers Can (Not) Predict

### v2

@@ Methods
The proof of Claim 2.6 builds on so-called chained Bell inequalities [9], and uses a result established by Barrett, Kent, and Pironio [4].
@@
We start by noting that the system Alice and Bob is assumed to be a bi-partite quantum system represented by a density operator $\rho\in\mathcal{S}(\mathcal{H}\otimes\mathcal{H})$ as the (mixed) state of the system, where $\mathcal{S}(\mathcal{H}\otimes\mathcal{H})$ is a convex subset of the self-adjoint trace-class operators on $\mathcal{H}=\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. We will only consider pure states of the form
|$\ket{\Psi}=\alpha\ket{00^{\prime}}+\sqrt{1-\alpha^{2}}\ket{11^{\prime}},$ (3.1)
@@
$E_{i}=\frac{1}{2}(\alpha_{0}^{i}\mathds{1}+\bm{\alpha}^{i}\cdot\bm{\sigma}),$ (3.3)
where $\bm{\alpha}^{i}=(\alpha_{x}^{i},\alpha_{y}^{i},\alpha_{z}^{i})$ a normalised vector, and $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ the Pauli matrices spanning the space of self-adjoint operators on $\mathbb{C}^{2}$.
@@
$\sigma_{n}:=\hat{n}\vec{\sigma}=\begin{pmatrix}\cos\vartheta&\sin\vartheta\\ \ \sin\vartheta&-\cos\vartheta\\ \end{pmatrix},$ (3.4)
@@
Suppose Maggie’s prediction box consists of a pointer yielding values for the variable $z$ for each run of the experiment. Suppose further that $z$ is distributed according to the measure $\mu(z)$.
$(RI)$ allows to define an objective probability distribution $p(x,y|a,b,\rho,z)$ which is well defined for all observers in the setup. Furthermore, (SD) ensures that all predictions of the three agents supervene on accessible variables and that all variables used are operationally manipulable. The quantum probabilities are reproduced if
@@
which we assume in the following. We have also made use of $(NC)$ which implies that $z$ is independent of Alice’s and Bob’s measurement settings. Without reliable intersubjectivity it wouldn’t be possible to ascribe a unique value to the settings $a,b$, and, in particular, the outcomes $x,y$ and Maggie’s variable $z$. For the values might be perspectival, i.e. depend on some other parameter in the experimental setup like the observers. It could also be that the measured outcomes don’t take on single values. As explained earlier, $(RI)$ is violated notably in many-worlds and relational interpretations. Moreover, reliability was also argued to be violated in some classes of hidden variable theories. There is no guarantee that Maggie’s outcomes are objectively defined for all observers, but importantly, there is also no guarantee that if they are, they are accessible to observers other than Maggie. For example, in pilot wave theory the configurations describing Maggie depend on the total wavefunction of the setup, and thus on Alice’s measurement by virtue of the dynamical laws for the particle configurations. Interestingly, the theorem shows that for predictive advantage of any kind and all measurements settings of Alice a contradiction arises.
@@
For the special case of Bell’s experiment suppose the predictive advantage consists of deterministic predictions for all measurements scenario. Let $p(x,y|a,b,z)$ be the predictions of an empirical extension. A deterministic prediction entails that the outcomes $x,y$ are functions of the settings $a,b$ and extension variable $z$. That is, $(x,y)=(f(a,b,z),g(a,b,z))$, and so
$p(x,y|a,b,z)$ $=\delta_{(x,y),((f(a,b,z),g(a,b,z)))}$ (3.7)
@@
But when on top of that, the probabilities are to satisfy signal-locality, then
$p(x,y|a,b,z)$ $=p(x|a,b,z)p(y|a,b,z)$ (3.8)
@@
Thus, the predictions describe uncorrelated and signal-local correlations that neither depend on the outcomes of the other site nor its settings. But such cannot reproduce all quantum correlations due to Bell’s theorem. Therefore, $p(x,y|a,b,z)$ cannot be (fully) deterministic (cf. also [35], [31], [12], and [18]). Thus, we conclude that any empirical extension of quantum mechanics necessarily makes indeterministic predictions for at least some settings and outcomes. This presents a first step for the result of this work. but note that Bell’s theorem does not provide proof the empirical completeness of quantum theory. What we demanded from a theory with predictive advantage was the possibility of finding some decomposition of quantum probabilities. That is, it need not be deterministic in general, and, in particular, not necessarily deterministic for the outcomes of all measurements. In fact, for a Bell setup with two measurement settings on each site probabilistic theories with predictive advantage can be constructed.[21]
@@
$I_{N}:=$ $\langle|x_{\vartheta_{1}}-y_{\vartheta_{1}}|\rangle+\langle|y_{\vartheta_{1}}-x_{\vartheta_{2}}|\rangle+\langle|x_{\vartheta_{2}}-y_{\vartheta_{2}}|\rangle$ (3.9)
@@
where the expectation value is defined as $\langle A\rangle:=\sum_{i=0}^{1}ip(A=i)=p(A=1)$. By defining $x_{\vartheta_{N+1}}:=x_{\vartheta_{1}}+1~{}(\bmod~{}2)$ it can be written as $I_{N}=\sum\limits_{n=1}^{N}(\langle|x_{\vartheta_{n}}-y_{\vartheta_{n}}|\rangle+\langle|y_{\vartheta_{n}}-x_{\vartheta_{n+1}}|\rangle)$. Up to minor details this expression corresponds to the so-called chained Bell inequalities first introduced by [9]. See in particular the generalisation to arbitrary dimensions in [4] who we closely follow in the subsequent presentation. This is the generalised Bell inequality. [^14] For $N=2$, $I_{N}\geq 1$ coincides with the familiar CHSH inequality.
@@
$\langle|x_{\vartheta_{k}}-y_{\vartheta_{l}}|\rangle$ $=p(|x_{\vartheta_{k}}-y_{\vartheta_{l}}|=1)$ (3.10)
@@
This holds for all terms but the last, where the argument is $1$ if and only if both outcomes are equal. For the latter case, the relevant probability is thus
[…]
Summing up gives the generalised Bell measure for the entangled state 3.1 with $N$ arbitrary measurement settings for Alice and Bob, i.e.
$I_{N}=\sum\limits_{n=1}^{N}(\langle|x_{\vartheta_{n}}-y_{\vartheta_{n}}|\rangle+\langle|y_{\vartheta_{n}}-x_{\vartheta_{n+1}}|\rangle)\\ =\sum_{n=1}^{N}\left[\sin^{2}(\frac{\vartheta_{n}}{2}-\frac{\vartheta_{n}^{\prime}}{2})+\sin^{2}(\frac{\vartheta^{\prime}_{n}}{2}-\frac{\vartheta_{n+1}}{2})\right]\\ -(\alpha\sqrt{1-\alpha^{2}}-\frac{1}{2})\sum_{n=1}^{N}\left[\sin\vartheta_{n}\sin\vartheta_{n}^{\prime}+\sin\vartheta_{n}^{\prime}\sin\vartheta_{n+1}\right].$
Interestingly, $I_{N}$ can be related to the marginal probabilities for Alice’s outcomes, and, as a result, to the hypothetical predictions of Maggie. This crucial insight was observed by [4]. Barrett et al.’s result requires two assumptions. First, the probabilities are no-signalling—which was introduced as signal-locality $(SL)$ here)—, and setting independence—the assumption of no-conspiracy $(NC)$. Thus, we can apply their result to the probability assignment of Alice’s outcomes conditioned on Maggie’s variable.
@@
In the quantum case the difference would vanish for maximally entangled states if Maggie didn’t have any predictive advantage. For Alice’s outcomes would be equally likely. Thus, on the assumption of predictive advantage $(PA)$ it follows that the difference cannot vanish. That is, for all predictions of Maggie there exists an $\varepsilon\in\mathbb{R}$ such that
$|p(x|a=\vartheta_{k},z)-p(1-x|a={\vartheta_{k}},z)|>\varepsilon.$ (3.12)
@@
$\min\limits_{N,\vartheta_{1},...,\vartheta_{N},\vartheta^{\prime}_{1},...,\vartheta_{N}}I_{N}.$ (3.14)
Equation 3.1 contains a term that only depends on the difference of measurement angles and is independent of the parameter $\alpha$. The second term vanishes when the state is maximally entangled, i.e. $\alpha=\frac{1}{\sqrt{2}}$. In this case one can then indeed find a set of measurements for which the generalised Bell measure contradicts the existence of predictive advantage, i.e. Equation 3.12.
@@
$\vartheta_{l}:=\frac{\pi}{N}(l-\frac{1}{2}),l=1,...,N.$ (3.16)
@@
$I_{N}\leq 2N\left(\frac{\pi}{4N}\right)^{2}\leq\frac{\varepsilon}{2}<\varepsilon$ (3.17)
@@
Conclusions
Can post-quantum theories have predictive advantage whilst conforming to the Born rule on average? We showed that unless epistemic boundaries exist between observers, the answer is no: quantum uncertainty is a fundamental limit to Nature’s predictability. For any hypothetical predictive advantage there exists a quantum measurement scenario with a finite number of settings which implies that predictive advantage would violate signal-locality. Giving up the assumption of ‘reliable intersubjectivity’ does in principle enable (subjective) predictive advantage (and even knowledge about the outcomes of non-commuting observables), but implies an epistemic boundary between different quantum observers.
To our knowledge this is the first rigorous result on predictive completeness addressing locality and the mutual epistemic reality of macroscopic observers.
We left open the question whether (operational) principles—other than signal-locality—can be invoked for a claim about empirical completeness. For example, these could be dynamical or methodological assumptions about the prediction improving variables. Moreover, no general proof is currently available for arbitrary single systems. The impossibility of predictive advantage for single systems may be established on grounds of non-contextuality—as a generalisation of locality.
In sum, if we assume that observers can reliably communicate and can’t send superluminal signals, the contemporary mainstream view that quantum predictions are irreducibly probabilistic becomes a precise theorem. Consequently, not only can quantum mechanical measurement outcomes not be predicted with certainty, the signal-local predictions of any theory must be exactly the quantum probabilities.
Acknowledgments

### v3

@@ Appendix A Methods
The proof of Claim 8 builds on so-called chained Bell inequalities [42], and uses a result established by Barrett, Kent, and Pironio [6].
@@
We start by noting that the system Alice and Bob share is assumed to be a bi-partite quantum system represented by a density operator $\rho\in\mathcal{S}(\mathcal{H}\otimes\mathcal{H})$ as the (mixed) state of the system, where $\mathcal{S}(\mathcal{H}\otimes\mathcal{H})$ is a convex subset of the self-adjoint trace-class operators on $\mathcal{H}=\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. We will only consider pure states of the form
|$\ket{\Psi}=\alpha\ket{00^{\prime}}+\sqrt{1-\alpha^{2}}\ket{11^{\prime}},$ (9)
@@
$E_{i}=\frac{1}{2}(\alpha_{0}^{i}\mathds{1}+\bm{\alpha}^{i}\cdot\bm{\sigma}),$ (11)
where $\alpha_{0}^{i}$ a real number, $\bm{\alpha}^{i}=(\alpha_{x}^{i},\alpha_{y}^{i},\alpha_{z}^{i})$ a normalised vector, and $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ the Pauli matrices spanning the space of self-adjoint operators on $\mathbb{C}^{2}$.
@@
$\sigma_{n}:=\hat{n}\cdot\vec{\sigma}=\begin{pmatrix}\cos\vartheta&\sin\vartheta\\ \ \sin\vartheta&-\cos\vartheta\\ \end{pmatrix},$ (12)
@@
Suppose Maggie’s prediction box consists of a pointer yielding values for the variable $z$ for each run of the experiment. Suppose further that $z$ is distributed according to the measure $\mu(z|a,b,\rho)$, which generally depends on the measurement settings $a,b$ and the prepared quantum state $\rho$.
$(RI)$ allows to define an objective probability distribution $p(x,y|a,b,\rho,z)$ which is well-defined for all observers in the setup. Furthermore, (SD) ensures that all predictions of the three agents supervene on accessible variables and that all variables used are operationally manipulable. The quantum probabilities are reproduced if
@@
which we assume in the following. We have also made use of $(NC)$ which implies that $z$ is independent of Alice’s and Bob’s measurement settings. Without reliable intersubjectivity it wouldn’t be possible to ascribe a unique value to the settings $a,b$, and, in particular, the outcomes $x,y$ and Maggie’s variable $z$. The values might be perspectival, i.e. depend on some other parameter in the experimental setup like the observers. It could also be that the measured outcomes don’t take on single values. As explained earlier, $(RI)$ is violated notably in many worlds and relational interpretations. Moreover, reliability was also argued to be violated in some classes of hidden variable theories. There is no guarantee that Maggie’s outcomes are objectively defined for all observers, but importantly, there is also no guarantee that if they are, they are accessible to observers other than Maggie. For example, in pilot wave theory the configurations describing Maggie depend on the total wavefunction of the setup, and thus on Alice’s measurement by virtue of the dynamical laws for the particle configurations. Interestingly, the theorem shows that for predictive advantage of any kind and all measurements settings of Alice a contradiction arises.
@@
For the special case of Bell’s experiment suppose the predictive advantage consists of deterministic predictions for all measurements. Let $p(x,y|a,b,z)$ be the predictions of an empirical extension. A deterministic prediction entails that the outcomes $x,y$ are functions of the settings $a,b$ and extension variable $z$. That is, $(x,y)=(f(a,b,z),g(a,b,z))$, and so
$p(x,y|a,b,z)$ $=\delta_{(x,y),((f(a,b,z),g(a,b,z)))}$ (15)
@@
But when on top of that, the probabilities are to satisfy no-signalling, then
$p(x,y|a,b,z)$ $=p(x|a,b,z)p(y|a,b,z)$ (16)
@@
Thus, the predictions describe uncorrelated and no-signalling correlations that neither depend on the outcomes of the other site nor its settings. However, such cannot reproduce all quantum correlations due to Bell’s theorem. Therefore, $p(x,y|a,b,z)$ cannot be (fully) deterministic (cf. also [43], [44], [45], and [1]). Thus, we conclude that any empirical extension of quantum mechanics necessarily makes indeterministic predictions for at least some settings and outcomes. This presents a first step for the result of this work, but note that Bell’s theorem does not provide proof the empirical completeness of quantum theory. What we demanded from a theory with predictive advantage was the possibility of finding some decomposition of quantum probabilities. That is, it need not be deterministic in general, and, in particular, not necessarily deterministic for the outcomes of all measurements. In fact, for a Bell setup with two measurement settings on each site probabilistic theories with predictive advantage can be constructed.[35]
@@
$I_{N}:=$ $\langle|x_{\vartheta_{1}}-y_{\vartheta_{1}}|\rangle+\langle|y_{\vartheta_{1}}-x_{\vartheta_{2}}|\rangle+\langle|x_{\vartheta_{2}}-y_{\vartheta_{2}}|\rangle$ (17)
@@
[…]
@@
$\langle|x_{\vartheta_{k}}-y_{\vartheta_{l}}|\rangle$ $=p(|x_{\vartheta_{k}}-y_{\vartheta_{l}}|=1)$ (18)
@@
This holds for all terms but the last, where the argument is $1$ if and only if both outcomes are equal. For the latter case, the relevant probability is thus {dmath} ⟨|x_ϑ_k-y_ϑ_l-1|⟩=p(|x_ϑ_k-y_ϑ_l-1|=1)
=tr(E_0^ϑ_k⊗E_0^ϑ_lρ)+tr(E_1^ϑ_k⊗E_1^ϑ_lρ).
Summing up gives the generalised Bell measure for the entangled state 9 with $N$ arbitrary measurement settings for Alice and Bob, i.e. {dmath} I_N=∑_n=1^N (⟨|x_ϑ_n-y_ϑ_n|⟩+⟨|y_ϑ_n-x_ϑ_n+1|⟩)
=∑_n=1^N[sin^2(ϑn2-ϑn’2)+sin^2(ϑn2-ϑn+12)]
-(α1-α^2-12)∑_n=1^N[sinϑ_nsinϑ_n’+sinϑ_n’sinϑ_n+1].
Interestingly, $I_{N}$ can be related to the marginal probabilities for Alice’s outcomes, and, as a result, to the hypothetical predictions of Maggie. This crucial insight was observed by [6]. Barrett et al.’s result requires two assumptions. First, the probabilities are no-signalling—which was introduced as no-signalling $(SL)$ here)—, and setting independence—the assumption of no-conspiracy $(NC)$. Thus, we can apply their result to the probability assignment of Alice’s outcomes conditioned on Maggie’s variable.
@@
In the quantum case, the difference would vanish for maximally entangled states if Maggie didn’t have any predictive advantage. For Alice’s outcomes would be equally likely. Thus, on the assumption of predictive advantage $(PA)$ it follows that the difference cannot vanish. That is, for all predictions of Maggie there exists an $\varepsilon\in\mathbb{R}$ such that
$|p(x|a=\vartheta_{k},z)-p(1-x|a={\vartheta_{k}},z)|>\varepsilon.$ (20)
@@
$\min\limits_{N,\vartheta_{1},...,\vartheta_{N},\vartheta^{\prime}_{1},...,\vartheta_{N}}I_{N}.$ (22)
Equation A contains a term that only depends on the difference of measurement angles and is independent of the parameter $\alpha$. The second term vanishes when the state is maximally entangled, i.e. $\alpha=\frac{1}{\sqrt{2}}$. In this case, one can then indeed find a set of measurements for which the generalised Bell measure contradicts the existence of predictive advantage, i.e. Equation 20.
@@
$\vartheta_{l}:=\frac{\pi}{N}(l-\frac{1}{2}),l=1,...,N.$ (24)
@@
$I_{N}\leq 2N\left(\frac{\pi}{4N}\right)^{2}\leq\frac{\varepsilon}{2}<\varepsilon$ (25)